# Molecular clouds: do they deserve a non-Gaussian description?

### v1

@@ 6 Comparing pairs and datasets
The previous section confronted multiple sets of statistics to assess their information content through their level of degeneracy. In this section, we fix the set of statistics: $\phi=\phi_{\text{final}}$, composed of 17 coefficients (7 log-Gaussian descriptors and 10 RWST statistics), and use it to define a distance between maps. We illustrate this distance by evidencing closest pairs of images in a dataset, as well as between different datasets such as observations and simulations.
@@ 6.1 Defining a morphological distance
Our objective is to define a distance between two patches based on $\phi_{\text{final}}$. One of our requirements is to enable a comparison between distance values for different pairs, so that these pairs can be ordered according to the morphological proximity of their patches. This requirement prevents us from using the statistical compatibility diagnostic introduced earlier. Indeed, since it is weighted by the local variance of each patch’s statistics, this can lead, for example, to some patches in a dataset being compatible with almost all the others simply because their spatial variance is very high. This could also encourage a simulation to approach an observation, following this criterion, by arbitrarily increasing its variance $\operatorname{Var}\phi(x_{SIM})$ without focusing on reducing the discrepancy of its average statistical properties $\hat{\phi}(x_{SIM})-\hat{\phi}(x_{OBS})$. This property was purposely used to act as a penalization when confronting different sets of statistics in the previous section, but is no longer desired for pairs’ ordering, once $\phi$ is fixed.
To build a distance that avoids this drawback, we choose instead to normalize it by the variability of the statistics evaluated over the entire observational dataset. We thus use the following distance:
@@
which is normalized by the spanning of the estimated values $\hat{\mu}_{i}$ of $\phi_{\text{final}}$ over all maps of a given dataset $\mathcal{D}$:
@@
where the brackets indicate an average over $\mathcal{D}$. Note, however, that it is difficult to interpret the value of $d^{2}_{\mathcal{D}}$ in absolute terms. Indeed, the $M_{\mathcal{D}}$ term does not describe a typical variance for a given process, but describes the variety of morphologies encountered in the entire dataset, that can be wide, as for the MC data investigated here. Unlike statistical compatibility diagnostics, the $d^{2}_{\mathcal{D}}$ distance is therefore modified by the addition or removal of maps in the $\mathcal{D}$ dataset, and can be affected by the presence of outliers.
In the following, we work with different datasets. For instance, we aim at comparing the minimum distance between observations and simulations to the typical distance of the closest pairs of observations. To do so, we use in this paper a common metric $M_{\text{obs}}$ computed on the observation dataset, as well as the associated $d^{2}_{\text{obs}}$ distance.
@@ 6.2 Closest pairs
We use the $d^{2}_{\text{obs}}$ distance to identify the closest pairs of patches that can be found between two datasets. To do so, we report in Fig. 15 the cumulative distributions of distances associated to pairs of observation patches (in gray), as well as to pairs made of one observation patch and one logFBM, DTD, and simulation patch (in red, purple, and blue, respectively). For each case, we pick up six of the typical closest pairs [^9] We do not investigate the $\sim 1\%$ closest pairs of observations, since the comparison between observations is slightly biased with respect to the other ones. The reason is that, in that case only, we have pairs made of non independent patches, such as neighboring patches in the sky or two slightly different scaled versions of the same region. This corresponds to approximately $1\%$ of the pairs, that we therefore exclude., as well as six relatively distant pairs, that we show in Fig. 16. In dashed lines with the same colors, we also report the cumulative distributions of distances associated to pairs of patches from a common dataset (logFBM, DTD [^10] Note that the $\sim 10$ closest pairs of DTD images, located at $d^{2}_{\text{obs}}\lessapprox 1$ are artifacts in this dataset corresponding in practice to almost identical images., and simulation patches, respectively).
Figure 15: Cumulative distributions of $d^{2}_{\text{obs}}$ distances between pairs extracted from multiple datasets. The different curves correspond to different choices of datasets from which the two patches of a given pair are extracted. This distance is based on $\phi_{\text{final}}$, a set of 17 coefficients (7 log-Gaussian descriptors and 10 RWST statistics). The same metric $M_{\text{obs}}$ is used for all distances and is defined based on the observational dataset. The stars in the bottom sub-panel correspond to ”close” and ”distant” pairs shown in Fig. 16. An horizontal line at the value $1\%$ is drawn to represent the proportion of pairs of observations that are neighboring patches in the sky. For these pairs, the distance is underestimated and thus the gray curve below this line is not meaningful. Same applies for pairs of simulations (blue dashed curve).
@@
Closest (OBS, OBS) pairs and (OBS, logFBM) pairs are visually rather similar, while distant pairs look very different. This relative agreement between morphological proximity as probed by our statistical distance as well as by human vision is encouraging, because it should be satisfied by an ideal distance. However, it remains far from being an exhaustive diagnostic.
The closest (OBS, logFBM) pairs have a distance $d^{2}_{\text{obs}}\sim 2$, which marks a high agreement. Indeed, only $\sim 0.5\%$ of (OBS, OBS) or (logFBM, logFBM) pairs exhibit smaller distance values, and we have already mentioned that the closest 1% of (OBS, OBS) pairs are very similar by construction. Comparatively, simulations are more distant to observations: the closest pairs between observations and simulations have a distance $d^{2}_{\text{obs}}\sim 8$, four times larger than the closest (OBS, logFBM) pairs, when already $\sim 10\%$ of pairs of observations exhibit smaller distances (see Sec. 6.3 for more details on this interpretation). As expected, these closest (OBS, SIM) pairs are visually less similar than the closest (OBS, OBS) or (OBS, logFBM) pairs (left panels of Fig. 16), but are still not so different, for instance with respect to the more distant pairs, that can be seen in the right panels of Fig. 16.
@@
The closest (OBS, logFBM) samples obtained here are very diffuse regions of MCs such as Corona Australis, Polaris Flare, and Ophiuchus. This is not surprising because for MCs, diffuse regions are closer to logFBM models than dense regions, whose PDFs are known to deviate from log-normality. However, these diffuse regions are still supposed to exhibit coherent structures that should induce deviations from logFBM models. Here, such deviations are found to be small. We note, however, that this proximity only means that certain diffuse regions are close to logFBM processes relative to the total variability of the observational dataset, and not necessarily that they are well described by such models, or could not be distinguished by the previous compatibility diagnostic. Moreover, this proximity to logFBM models is also partly due to contamination by the Cosmic Infrared Background (CIB), that has non negligible power at such low levels of column density, and is expected to Gaussianize the data, including their RWST statistics (Auclair et al., 2024). Incidentally, we remind that the observational dataset has some artifacts that make it deviate from an ”ideal” MC column density dataset, even if we try to limit as much as possible their effects (noise, finite resolution), as discussed in Sec. 4.
@@
On the contrary, the observations that are closest to the simulations of dense MCs correspond, unsurprisingly, to dense regions of MCs such as Ophiuchus, Aquila, and Serpens. Note that Ophiuchus exhibits patches that are close to these dense simulations, but also at least one patch that is close to a logFBM model, underlining the spatial variability of molecular cloud morphologies. This illustrates the difficulty to treat a MC as a single entity, and emphasizes the relevance of our local approach.
@@
The comparison between observations and DTD shows the limitations of our distance diagnostic. Indeed, the closest (OBS, DTD) pairs are found at a distance $d^{2}_{\text{obs}}\sim 2$ that is the typical distance between close (OBS, OBS) pairs or close (OBS, logFBM) pairs, although they are visually very different. This illustrates that, for a diagnostic based on a low-dimensional set of statistics, it is difficult to probe a distance over a family of processes that has such a wide variety of textures as DTD. On the contrary, because the simulations that are closest to observations are more distant ($d^{2}_{\text{obs}}\sim 8$), this suggests that the set of simulations does not intersect the set of observations and that $\phi_{\text{final}}$ is able to pinpoint this discrepancy, as discussed in the following subsection.
@@
Finally, it is quite impressive to see that we can build a distance diagnostic from a representation of dimension 17 only that still manages to identify morphological similarity between maps quite satisfactorily. This illustrates the possibility of constructing a highly informative but low-dimensional description tailored to a family of processes from an ensemble of compressed sets of usual statistics. It should be stressed, however, that this study remains partial, notably because it is based mainly on the observation of a few close pairs in our dataset. Yet, the confrontation diagnostics studied in Sec. 5 showed that degenerate counterexamples remain largely in the minority. In addition, we lack solid baselines since it is inherently difficult to quantify visual impressions of morphological proximity, although some work has been done in this direction (Peek & White, 2021).
@@ 6.3 Interpreting the minimal distance between observations and simulations
A last question we tackle is whether the relatively high value of the minimal distance $d^{2}_{\text{obs}}$ that we get between observations and simulations indeed indicates distinct statistical properties of these two sets. More precisely, we question whether the set of observations’ moments $\{\bar{\mu}_{i}\}_{i\in\text{obs}}$ overlaps that of simulations $\{\bar{\mu}_{j}\}_{j\in\text{sim}}$, where $\bar{\mu}_{i}\equiv\mathbb{E}[\phi_{\text{final}}(x_{i})]$ stands for the expected value over a given process $i$. In practice, we aim at retrieving the minimal distance
between $\bar{\mu}_{i}$ and $\bar{\mu}_{j}$ over (OBS, SIM) pairs $(i,j)$. However, the distance $d^{2}_{\text{obs}}$ introduced previously is a statistical estimator, so that in general
$d^{2}_{ij}\neq\bar{d}^{2}_{ij},$
and in particular: $\mathbb{E}[d^{2}_{ij}]>\bar{d}^{2}_{ij}$. Indeed, the variance of the $\hat{\mu}_{i}$ estimator biases $d^{2}_{ij}$ with respect to $\bar{d}^{2}_{ij}$:
$\mathbb{E}[d^{2}_{ij}]\equiv\bar{d}^{2}_{ij}+b_{ij}.$ (18)
This non-negative bias, which boils down to:
prevents us from interpreting directly the value of $d^{2}_{ij}$ as $\bar{d}^{2}_{ij}$. Furthermore, its value increases with the amplitude of the fluctuations of $\hat{\mu}_{i}-\hat{\mu}_{j}$, and can thus change depending on the pair $(i,j)$ considered. This dependency on the pair also prevents us, without further check, from actually comparing $\bar{d}^{2}$ between different pairs based on the estimations $d^{2}$.
  • find an observation $i$ and a simulation $j$ that minimize the biased estimation $d^{2}_{ij}$. Such pairs are already reported in Fig.16 and are good candidates to minimize $\bar{d}^{2}_{ij}$. The goal then becomes to compare $\bar{d}^{2}_{ij}$ to the minimal value $\bar{d}^{2}_{ii^{\prime}}$, for $i^{\prime}$ another observation, independent from $i$. But $\bar{d}^{2}$ is unknown. Instead:
  • If, in addition the observation patch $x_{i^{\prime}}$ is such that spatial fluctuations of $\phi_{\text{final}}$ over its sub-patches are at least of the order of those estimated from the sub-patches of the simulation patch $x_{j}$, then based on Eq. 19, we have $b_{ii^{\prime}}\gtrapprox b_{ij}$. This allows then to use $d^{2}_{ij}-d^{2}_{ii^{\prime}}$ as an estimated lower bound for $\bar{d}^{2}_{ij}-\bar{d}^{2}_{ii^{\prime}}$.
We report in Fig. 17 such independent [^11] Indeed, only one of these pairs is made of patches retrieved from the same MC (Ophiuchus), but these patches do not overlap. pairs $(i,i^{\prime})$. In the first row, we fix $i$ associated to the patch of Ophiuchus $x_{i}$ that is the closest observation to simulations ($d^{2}_{ij}=7$, Fig.16). For that choice of patch, the reported distances $d^{2}_{ii^{\prime}}$ of neighboring observations $i^{\prime}$ are rather similar to $d^{2}_{ij}$. However, in the second row, where $i$ is associated to the patch of Aquila that is the second closest observation to simulations (as shown in Fig.16), we see other observations $i^{\prime}$, such as Serpens or Orion B, that are closer than the closest simulation: $d^{2}_{ii^{\prime}}=3<d^{2}_{ij}=8$. We checked that the fluctuations of $i^{\prime}$ are comparable to the one of $j$ so that $b_{ii^{\prime}}\gtrapprox b_{ij}$. This implies then $\bar{d}^{2}_{ij}-\bar{d}^{2}_{ii^{\prime}}\gtrapprox 5$ while $\bar{d}^{2}_{ii^{\prime}}\leq d^{2}_{ii^{\prime}}=3$. Hence, $\bar{d}^{2}_{ij}/\bar{d}^{2}_{ii^{\prime}}\gtrapprox 8/3$. This second case shows that one of the closest observation/simulation pair is definitely further apart than this specific observation is with other observations, at least according to $\phi_{\text{final}}$. This example also supports that the potential bias on the $d^{2}_{\text{obs}}$ estimators remains moderate for the study of (OBS, SIM) pairs ($\bar{d}^{2}_{ij}\gtrapprox 5$ and $d^{2}_{ij}=8$ implies $b_{ij}\lessapprox 3$).
@@
We conclude that the minimum (OBS, SIM) distance value, obtained on all possible pairs between those datasets, evidences a meaningful but moderated distinction between these datasets. Note that the $\phi_{\text{final}}$ statistics set may still have some degeneracies, particularly for the (OBS, SIM) comparison, and that accounting for these should likely deepen this gap. We however believe that current analysis illustrates the usefulness of such a distance, and leave a more detailed study for future work.
Figure 17: Comparing closest (OBS, SIM) pairs $(i,j)$ (gray boxes) to neighboring (OBS, OBS) pairs $(i,i^{\prime})$. Top row focuses on the closest observation to simulations, that is a patch of Ophiuchus, while bottom row focuses on the second closest patch, that is in Aquila cloud. In the latter case, neighboring (but independent) observations of Aquila are closer than the closest simulations. The color-bars can change from a pair to another.
@@ 7 Conclusions
In this paper, we aim at studying the diversity of morphologies of observed molecular clouds. To do so we construct a set of $\sim 500$ patches of size $512\times 512$ pixels, extracted at different resolutions from column density maps of 14 nearby clouds derived from the HGBS dust emission observations. We compute several sets of statistics (mean, variance, quantile function, power spectrum and scattering transform) from these maps, and we compare their informative power. To do so, we introduce a new methodology (Fig. 6) that allows to confront, without any supervision, two sets of summary statistics on their respective abilities to detect statistical incompatibility between pairs of processes in a given dataset.
Applying this methodology to this set of observations, we find that Gaussian statistics have degeneracies for the observational dataset, some of which can be lifted by one-point statistics (Fig. 7). We then show that even if log-Gaussian statistics are much less degenerate, a compressed set of scattering statistics still succeeds to demonstrate further degeneracies (Fig. 11). This confirms that the diversity of morphologies arising in these observed clouds cannot be sufficiently described by Gaussian nor log-Gaussian statistics. This means that using such descriptions, typically to compare numerical simulations with observational data, can lead to misleading projections: in addition to missing potential absolute discrepancies, observations with different properties can still be matched to the same simulation.
We apply the same diagnostic to simulations and to the set of logFBMs (Fig. 14), and find deviations from observations’ geometry. In particular, the outcome of this diagnostic strikingly supports the sufficiency of log-Gaussian statistics to discriminate between logFBM patches. Regarding observations and simulations, there remain deviations between their respective geometries, even though great care has been taken in this work to design robust and low dimensional sets of summary statistics. This supports the difficulty of transferring simulation-based priors [^12] Such as the outcome of a Fisher analysis, or a trained neural network. to observations, especially high-order information content, and also buttresses the supervision-detached approach developed in this paper, along with the choice to work with compressed and robust summary statistics.
From these results, we introduce a morphological distance $d^{2}_{\text{obs}}$ based on a set of summary statistics $\phi_{\text{final}}$ composed of 7 log-Gaussian and 10 RWST coefficients. The similarity probed by this distance is in agreement with visual impression when comparing observations with themselves, with logFBMs and with simulations (Fig. 16). It remains however insufficient to operate on datasets made of a wider diversity of textures such as DTD.
@@
  • the methodology we developed to confront summary statistics requires very few assumptions: it can operate in an unsupervised and very low data regime. Hence such methodology can easily be assimilated by the ISM community and applied on a wider set of statistics and physical tracers of data (velocities, polarization, temperature).
  • The low dimensional and analytical morphological embedding $\phi_{\text{final}}$ developed here allows for a directly interpretable comparison, for instance between different observed clouds, between observations and simulations or statistical models, but it also paves the way for the use of more sophisticated unsupervised learning techniques.
  • Leveraging saliency maps $\nabla_{\text{pixels}}d_{\text{obs}}^{2}(x_{i},x_{j})$, the distance $d^{2}_{\text{obs}}$ can be used to highlight the main areas responsible for morphological discrepancies between two patches $(x_{i},x_{j})$, broadening the scope of the work initiated by Peek & Burkhart (2019) to the unsupervised world of observations.
  • The confrontation methodology can be used to make a feature selection algorithm, in the spirit of the FRAME model developed by Zhu et al. (1998), but designed to optimize the comparison task of a non-parametric collection of processes $\{p_{i}\}_{i}$, instead of modeling a single process.
@@
  • reduce the uncertainty in the compatibility diagnostic $d^{2}_{\phi}$ due to the precision matrix estimation. A promising idea is to use a maximum entropy model conditioned on the data on which to perform the precision estimation.
  • Reduce the overall dependency of the distance $d^{2}_{\mathcal{D}}$ on the dataset $\mathcal{D}$, and in particular on its outliers. This can be done by building a localized metric, based on local estimations of the geometry, restricting $\mathcal{D}$ to $\mathcal{D}_{\text{loc}}[i]$, the neighbors of a given process $i$.

### v2

@@ 6 Comparing pairs and datasets
In the previous section, we confronted multiple sets of statistics to assess their information content through their level of degeneracy. In this section, we fixed the set of statistics: $\phi=\phi_{\text{final}}$, composed of 17 coefficients (seven log-Gaussian descriptors and ten RWST statistics), and used it to define a morphological distance between maps. We illustrate this distance by exhibiting closest pairs of images in a dataset, as well as between different datasets such as observations and simulations.
Figure 11: Cumulative distributions of $d^{2}_{\text{obs}}$ morphological distances between pairs extracted from multiple datasets. The different curves correspond to different choices of datasets from which the two patches of a given pair are extracted. This distance is based on $\phi_{\text{final}}$, a set of 17 coefficients (seven log-Gaussian descriptors and ten RWST statistics). The same metric $M_{\text{obs}}$ is used for all distances and is defined based on the observational dataset. The stars in the bottom sub-panel correspond to ”close” and ”distant” pairs shown in Fig. 12. An horizontal line at the value $1\%$ is drawn to represent the proportion of pairs of neighboring patches in the sky. These pairs made of patches with strong statistical dependence are likely to lead to an underestimated distance with respect to the other pairs. While these corrupted pairs might significantly affect the gray curve below this horizontal line, they can only have a negligible influence above that line. Same applies for pairs of simulations (blue dashed curve).
@@ 6.1 Defining a morphological distance
Our objective is to define a morphological distance between two patches based on $\phi_{\text{final}}$. One of our requirements is to enable a comparison between distance values for different pairs, so that these pairs can be ordered according to the morphological proximity of their patches. This requirement prevents us from using the statistical compatibility diagnostic introduced earlier. Indeed, since it is weighted by the local variance of each patch’s statistics, this can lead, for example, to some patches in a dataset being compatible with almost all the others simply because their spatial variance is very high. This could also encourage a simulation to approach an observation, following this criterion, by arbitrarily increasing its variance $\operatorname{Var}\phi(x_{\text{SIM}})$ without focusing on reducing the discrepancy of its average statistical properties $\hat{\phi}(x_{\text{SIM}})-\hat{\phi}(x_{\text{OBS}})$. This property was purposely used to act as a penalization when confronting different sets of statistics in the previous section, but is no longer desired for pairs’ ordering, once $\phi$ is fixed.
To build a morphological distance that avoids this drawback, we choose instead to normalize it by the variability of the statistics evaluated over the entire observational dataset. We thus use the following distance:
@@
which is normalized by the variance of $\hat{\mu}_{i}$ over all maps of a given dataset $\mathcal{D}$:
@@
where the brackets indicate an average over $\mathcal{D}$. We note, however, that it is difficult to interpret the value of $d^{2}_{\mathcal{D}}$ in absolute terms. Indeed, the $M_{\mathcal{D}}$ term does not describe a typical variance for a given process, but describes the variety of morphologies encountered accross the entire dataset, that can be broad, as seen in the MC data investigated here. Unlike statistical compatibility diagnostics, the $d^{2}_{\mathcal{D}}$ distance is therefore modified by the addition or removal of maps in the dataset $\mathcal{D}$, and can be affected by the presence of outliers.
In the following, we work with different datasets. For instance, we aim at comparing the minimum distance between observations and simulations to the typical distance of the closest pairs of observations. To do so, we use in this paper a unique metric, $M_{\text{obs}}$, computed on the observation dataset, as well as its associated $d^{2}_{\text{obs}}$ distance.
@@ 6.2 Closest pairs
We use the $d^{2}_{\text{obs}}$ morphological distance to identify the closest pairs of patches that can be found between two datasets. To do so, we report in Fig. 11 the cumulative distributions of distances associated with pairs of observation patches (in gray), as well as with pairs consisting of one observation patch and one logFBM, DTD, and simulation patch (in red, purple, and blue, respectively). For each case, we select six of the typical closest pairs [^12] We do not investigate the $\sim 1\%$ closest pairs of observations, since the comparison between observations is slightly biased with respect to the other ones. The reason is that, in that case only, we have pairs made of non-independent patches, such as neighboring patches in the sky or two slightly different scaled versions of the same region. This corresponds to approximately $1\%$ of the pairs, which we therefore exclude., as well as six relatively distant pairs, that we show in Fig. 12. In dashed lines with the same colors, we also report the cumulative distributions of distances associated to pairs of patches from a common dataset (logFBM, DTD [^13] We note that the $\sim 10$ closest pairs of DTD images, located at $d^{2}_{\text{obs}}\lessapprox 1$ are artifacts in this dataset corresponding in practice to almost identical images., and simulation patches, respectively).
@@
Closest (OBS, OBS) pairs and (OBS, logFBM) pairs are visually rather similar, while distant pairs look very different. This relative agreement between the morphological proximity, as probed by our statistical distance and by human vision, is encouraging, because it should be satisfied by an ideal distance. However, it remains far from being an exhaustive diagnostic.
[…]
Figure 13: Comparison of closest (OBS, SIM) pairs $(i,j)$ (gray boxes) to neighboring (OBS, OBS) pairs $(i,i^{\prime})$. The top row focuses on the closest observation to simulations, which corresponds to a patch of Ophiuchus, while the bottom row focuses on the second closest patch, which is in Aquila. In the latter case, neighboring (but independent) observations of Aquila are significantly closer than the closest simulations. The maps are shown in units of $10^{20}\text{cm}^{-2}$ but the colorbars can change from a pair to another.
@@
The closest (OBS, logFBM) samples obtained here are very diffuse regions of MCs such as Corona Australis, Polaris Flare, and Ophiuchus. This is not surprising because, for MCs, diffuse regions are closer to logFBM models than dense regions, whose PDFs are known to deviate from log-normality. However, these diffuse regions are still supposed to exhibit coherent structures that should induce deviations from logFBM models. Here, such deviations are found to be small. We note, however, that this proximity only means that certain diffuse regions are close to logFBM processes relative to the total variability of the observational dataset, and not necessarily that they are well described by such models, or could not be distinguished by the previous compatibility diagnostic. Moreover, this proximity to logFBM models is also partly due to contamination by the CIB which has non-negligible power at such low levels of column density, and is expected to Gaussianize the data, including their RWST statistics (Auclair et al., 2024). Incidentally, we remind that the observational dataset has some artifacts that make it deviate from an ”ideal” MC column density dataset, even if we try to limit as much as possible their effects (noise, finite resolution), as discussed in Sect. 4.
@@
On the contrary, the observations that are closest to the simulations of dense MCs correspond, unsurprisingly, to dense regions of MCs such as Ophiuchus, Aquila, and Serpens. We note that Ophiuchus exhibits patches that are close to these dense simulations, but also at least one patch that is close to a logFBM model, underlining the spatial variability of molecular cloud morphologies. This illustrates the difficulty to treat a MC as a single entity, and emphasizes the relevance of our local approach.
@@
The comparison between observations and DTD shows the limitations of our morphological distance diagnostic. Indeed, the closest (OBS, DTD) pairs are found at a distance $d^{2}_{\text{obs}}\sim 2$ that is the typical distance between close (OBS, OBS) pairs or close (OBS, logFBM) pairs, although they are visually very different. This illustrates that, for a diagnostic based on a low-dimensional set of statistics, it is difficult to probe a distance over a family of processes that has such a wide variety of textures as DTD. On the contrary, because the simulations closest to the observations are more distant ($d^{2}_{\text{obs}}\sim 8$), this suggests that the set of simulations does not intersect the set of observations and that $\phi_{\text{final}}$ is able to pinpoint this discrepancy, as discussed in the following subsection.
@@
Finally, it is quite impressive to see that we can build a distance diagnostic from a representation of dimension 17 only that still manages to identify morphological similarity between maps quite satisfactorily. This illustrates the possibility of constructing a highly informative but low-dimensional description tailored to a family of processes from an ensemble of compressed sets of typical statistics. It should be stressed, however, that this study remains partial, notably because it is based mainly on the observation of a few close pairs in our dataset. Yet, the confrontation diagnostics studied in Sect. 5 showed that degenerate counterexamples remain largely in the minority. In addition, we lack solid baselines since it is inherently difficult to quantify visual impressions of morphological proximity, although some work has been done in this direction (Peek & White, 2021).
@@ 6.3 Interpreting the minimal distance between observations and simulations
A last question we tackle is whether the relatively high value of the minimal distance $d^{2}_{\text{obs}}$ that we get between observations and simulations indeed indicates that these two sets of processes do not overlap. To do so, we select the observations that are closest to the simulations. As shown in Fig. 12, these correspond to patches of Ophiuchus and Aquila. Then, we report in Fig. 13 the closest observations to these patches, and sort them by proximity, according to $d^{2}_{\text{obs}}$.
  • Concerning the Ophiuchus patch (top row), only three observations patches seem to be closer ($d^{2}_{\text{obs}}\in\{6.5,7.2,7.3\}$) to it than the closest simulation ($d^{2}_{\text{obs}}=7.3$), and the difference in $d^{2}_{\text{obs}}$ is small.
  • However, concerning the Aquila patch (bottom row), there are numerous observations closer to it than the closest simulation ($d^{2}_{\text{obs}}=8$), and their distance is much smaller (down to $d^{2}_{\text{obs}}=3.1$).
@@
We further interpret this discrepancy between distances in Appendix F. We conclude that the minimum (OBS, SIM) distance value, obtained on all possible pairs between those datasets, evidences a meaningful but moderate distinction between these datasets. We note that the $\phi_{\text{final}}$ statistics set may still have some degeneracies, particularly for the (OBS, SIM) comparison, and that accounting for these should likely deepen this gap. However, we believe that the current analysis illustrates the usefulness of such a morphological distance, and leave a more detailed study for future work.
@@ 7 Conclusions
In this paper, we aim to study the diversity of morphologies of observed molecular clouds. To do so, we construct a set of $\sim 500$ patches of size $512\times 512$ pixels, extracted at different resolutions from column density maps of 14 nearby clouds derived from the HGBS dust emission observations. We compute several sets of statistics (mean, variance, quantile function, power spectrum and scattering transform) from these maps, and we compare their informative power. To do so, we introduce a new methodology (Fig. 5) that allows us to confront, without any supervision, two sets of summary statistics on their respective abilities to detect statistical incompatibility between pairs of stochastic processes of a given family.
Applying this methodology to this set of observations, we find that Gaussian statistics have degeneracies for the observational dataset, some of which can be lifted by one-point statistics (Fig. 6). We then show that even if log-Gaussian statistics are much less degenerate, a compressed set of scattering statistics still succeeds to demonstrate further degeneracies (Fig. 10 .b). This confirms that the diversity of morphologies arising in these observed clouds cannot be sufficiently described by either Gaussian or log-Gaussian statistics. This means that using such descriptions, typically to compare numerical simulations with observational data, can lead to misleading conclusions: in addition to missing potential absolute discrepancies, observations with different statistical properties can still be matched to the same simulation.
We apply the same diagnostic to simulations (Fig. 10.c) and to the set of logFBMs (Fig. 10 .a), and find deviations from the statistical structure of the set of observations. In particular, the outcome of this diagnostic strikingly supports the sufficiency of log-Gaussian statistics to discriminate between logFBM patches. Regarding observations and simulations, there remain deviations between their respective statistical structures, even though great care has been taken in this work to design robust and low-dimensional sets of summary statistics. This supports the difficulty of transferring statistical properties learned on simulations to observations, especially those properties relative to higher-order moments, as mentioned in the caveat n°2 in the introduction. This also buttresses the supervision-detached approach developed in this paper, along with the choice to work with compressed and robust summary statistics.
From these results, we introduce a morphological distance $d^{2}_{\text{obs}}$ based on a set of summary statistics $\phi_{\text{final}}$ composed of seven log-Gaussian and ten RWST coefficients. The similarity probed by this distance is in agreement with visual impression when comparing observations with themselves, with logFBMs and with simulations (Fig. 12). However, it remains insufficient to operate on datasets made of a wider diversity of textures such as DTD.
@@
  • The methodology we developed to confront summary statistics requires very few assumptions: it can operate in an unsupervised and very low data regime. Hence, such methodology can easily be assimilated by the ISM community and applied on a wider set of statistics and physical fields (velocities, polarization, temperature).
  • The low-dimensional and analytical morphological embedding $\phi_{\text{final}}$ developed here allows us for a statistically interpretable comparison, for instance between different observed clouds, between observations and simulations or statistical models. It also paves the way for the use of more sophisticated unsupervised learning techniques.
  • Leveraging saliency maps $\nabla_{\text{pixels}}d_{\text{obs}}^{2}(x_{i},x_{j})$, the morphological distance $d^{2}_{\text{obs}}$ can be used to highlight the main areas responsible for morphological discrepancies between two patches $(x_{i},x_{j})$, broadening the scope of the work initiated by Peek & Burkhart (2019) to the unsupervised world of observations.
  • The present work focuses on the statistical properties of molecular clouds. In particular, it assumes that patches with incompatible statistics should be considered as realizations of different processes. However, a potential caveat of this approach, when applied to nonstationary processes such as dense molecular clouds, could be to find statistical incompatibilities between patches which are not particularly relevant from a physical point of view. While we do not further discuss this issue in the present paper, we think it could deserve further work.
  • The confrontation methodology can be used to make a feature selection algorithm, in the spirit of the FRAME model developed by Zhu et al. (1998), but designed to optimize the comparison task of a nonparametric collection of processes $\{p_{i}\}_{i}$, instead of modeling a single process.
@@
  • Reduce the uncertainty in the compatibility diagnostic $d^{2}_{\phi}$ due to the precision matrix estimation. A promising idea is to use a maximum entropy model conditioned on the data  (Bruna & Mallat, 2019; Allys et al., 2019; Zhang & Mallat, 2021) on which to perform the precision estimation.
  • Reduce the overall dependency of the morphological distance $d^{2}_{\mathcal{D}}$ on the dataset $\mathcal{D}$, and in particular on its outliers. This can be done by building a localized metric, reducing the scope of the averages in Eq. 18 from $\mathcal{D}$ to $\mathcal{D}_{\text{loc}}[i]$, the set of neighbors of a given process $i$.